# Eco-evolutionary dynamics of adapting pathogens and host immunity

### v2

@@ Abstract
The traveling wave framework has been extensively used in this context as it allows for a straightforward way to approach the prediction problem: each variant is assumed to have a fixed fitness relative to other variants, and inferring the fitness of all competing variants should allow prediction of the future composition of the population. Indeed, current methods typically infer the instantaneous growth advantage of a strain based on past and present circulation and then project this growth advantage forward in time Luksza and Lässig (2014); Neher et al. (2014); Huddleston et al. (2020).
@@
In terms of modeling, a drawback of the traveling wave approach is the lack of explicit representation of the epidemiological dynamics and of the host’s immunity. Indeed, fitness is only an effective parameter that summarizes the complex interplay between viral antigenic properties and the hosts’ immune systems. As such, it cannot explicitly describe important phenomena such as the build-up of host immunity to new variants, variant specific immunity, or the interaction between strains through antigenic cross-reactivity. Taking the hosts’ immunity and viral cross-immunity into account has the potential to strongly improve predictions Meijers et al. (2023) or explain why prediction is difficult (Barrat-Charlaix et al., 2021).
@@
The interaction between epidemiological dynamics and hosts’ immunity are often modeled using generalizations of the Susceptible-Infected-Recovered model (SIR) to include multiple viral strains Gupta et al. (1998); Gog and Grenfell (2002). In this setting, the natural question is that of the ultimate fate of the pathogen: will it go extinct, diversify to the point of speciation, or reach the so-called Red Queen State where new strains continuously replace old ones Yan et al. (2019); Marchi et al. (2021); Chardès et al. (2023); Rouzine and Rozhnova (2018). To remain tractable, these studies typically approximate population immunity as a low-dimensional landscape in which the viral population evolves and ignore the complex heterogeneity in the immunity of different individuals. Furthermore, immunity is often assumed to be long lived and evolution of the pathogen in a stable low dimensional landscape gives rise to traveling waves.
@@
Here, we study how novel variants of a virus shape the host population’s immunity, which in turn changes their own growth dynamics. To do so, we use a multi-strain SIR model combining immune waning and heterogeneous immunity of the hosts. Such heterogeneity has been demonstrated for influenza virus in individuals of different ages Lee et al. (2019); Welsh et al. (2023). We show that this model generically leads to a situation where novel immune evasive variants emerge. In a homogeneous population of hosts, this leads to a succession of selective sweeps where novel variants compete against each other and replace previously circulating variants. However, in a heterogeneous population with more rapid waning of immunity, initially growing variants lose their selective advantage before reaching fixation due to immunological adjustment of the host population. The phenomenology of our model is reminiscent of ecological systems, where adaptation by one species shifts the global equilibrium and distribution of other species but does not necessarily result in a selective sweep. In these systems, adaptation can usually not be modeled by a scalar fitness parameter for each strain Tikhonov and Monasson (2018).
@@
Strain dynamics in our model differ qualitatively from what is expected in the traveling wave scenario. While adaptive mutations are highly overrepresented, they cease having a growth advantage when reaching intermediate frequencies, a process we call “expiring fitness”. Once the fitness effect of a mutation has expired, its frequency randomly changes up or down as subsequent adaptive mutations occur on the same or on different genomic backgrounds.
@@
We describe the interaction of several viral strains and host immunity using a Susceptible/Infected compartmental model, similar to those used in Gog and Grenfell (2002); Yan et al. (2019). In the most general form, the model describes $N$ variants of the virus labeled $a\in\{1\ldots N\}$ circulating among $M$ groups of hosts with distinct exposure histories labeled $i\in\{1\ldots M\}$ (immune groups). These groups could be different age cohorts or be geographically separated. For each group $i$, we define compartments $I_{i}^{a}$ and $S_{i}^{a}$ as respectively the number of individuals of this group infected or susceptible to strain $a$. We assume that the total population of hosts is $1$ so that we always have $0\leq I^{a}_{i},S^{a}_{i}\leq 1$, and values of $I^{a}_{i}$ and $S^{a}_{i}$ can be interpreted as fractions of the host population.
@@
In the rest of this work, we will often consider the case $C_{ij}=1/M$, corresponding to the case where hosts of different groups are equally likely to encounter members of any other group. Note that in this case, if we define the total number of susceptible or infected by $a$ as $S^{a}=\sum_{i}S_{i}^{a}$ and $I^{a}=\sum_{i}I_{i}^{a}$, we can sum over the immune groups in the above equation to obtain $\dot{I}^{a}=(\alpha S^{a}-\delta)I^{a}$. We consequently define the reproductive number of strain $a$ as $R^{a}=\alpha S^{a}/\delta$. $I^{a}$ grows when $R^{a}>1$ and declines when $R^{a}<1$. The equilibrium susceptibility is $S^{a}=\delta/\alpha$, such that $R^{a}=1$.
When a host of group $i$ is infected by strain $b$, it also has a chance $0\leq K_{i}^{ab}\leq 1$ to gain immunity to strain $a$. Thus, $S_{i}^{a}$ decreases at a rate proportional to $K_{i}^{ab}$ and to the number of hosts infected by $b$ for every strain $b$. Since susceptibles to $a$ are depleted by infections from other strains, the dynamics of all strains are coupled. This coupling is determined by the matrices $\mathbf{K}_{i}$ of dimension $N\times N$, which in general differ between groups $i$ with different prior exposure history. Additionally, waning of immunity at a rate $\gamma$ causes immune hosts to re-enter the susceptible compartment. We can now write the dynamics of $S_{i}^{a}$ as
@@
An important property of our model is that the probability to generate cross-immunity can differ between groups. The motivation is that strains $a$ and $b$ may be perceived as antigenically different by some immune systems, leading to a low $K_{i}^{ab}$, but as highly similar by others, leading to $K_{i}^{ab}\simeq 1$. Such a heterogeneous response by different immune systems has been observed experimentally in the case of influenza: in Lee et al. (2019); Welsh et al. (2023) for instance, it was found that a given mutation in an influenza strain may allow escape from the antibodies of some individuals, i.e. low $K_{i}^{ab}$, while it had little effect for the serum of other individuals, i.e. high $K_{i}^{ab}$. Heterogeneous immune response could be caused by varying histories of strain exposure for different individuals, for instance due to difference in age or geographical region. If immune groups correspond to age cohorts, mixing between groups is rapid and we can simplify the connectivity between groups to $C_{ij}=1/M$. If immune groups are shaped by geographic differences in exposure, the connectivity would be close to 1 on the diagonal ($1-C_{ii}\ll 1$) while off-diagonal terms would be small ($C_{ij}\ll 1$ for $i\neq j$).
@@
In the first version of this scenario, we will only consider one immune group, that is $M=1$. In this case, we can skip the indices $i,j\in\{1\ldots M\}$, and we only have one cross-immunity matrix $\mathbf{K}$ that we parametrize as
@@
$\begin{split}\dot{S}^{a}&=-\alpha S^{a}\sum_{b\in\{wt,v\}}K^{ab}I^{b}+\gamma(1-S^{a}),\\ \dot{I}^{a}&=\alpha S^{a}I^{a}-\delta I^{a}.\\ \end{split}$ (4)
We can immediately derive the equilibrium state for this simplified case. For the susceptibles we find $S^{a}=\delta/\alpha$ for all strains $a$, while the equilibrium prevalence is determined by the inverse of the cross-immunity matrix $\mathbf{K}$:
@@
In our invasion scenario, only the wild-type is initially present in the population, that is $I^{v}=0$ at $t<0$. The details of the equilibrium reached by the system in the absence of the mutant is given in the SI. Importantly, in this initial state, the number of hosts susceptible to the variant is always larger than the equilibrium value $\delta/\alpha$ unless $f=1$, but smaller than $1$ unless $f=0$. This comes from the fact that due to cross-immunity, part of the hosts are immune to the variant even if it is not yet circulating. As a result, when the variant is introduced at $t=0$, for instance because it emerged by mutation, its growth rate is initially positive.
Figure 1: Invasion scenario for Top row: one immune group, Middle row: $M=10$ immune groups and fast mixing $C_{ij}=1/M$ and Bottom row $M=10$ immune groups and slow mixing $C_{ij}=1/10M$. Other parameters are the same for all rows: in units of $\delta$, we set $\alpha=3$ and $\gamma=5\cdot 10^{-3}$, and $f=0.65$, $b=0.8$, $\varepsilon=0.99$. For both rows, graphs represent: Left: number of hosts infectious with the wild-type and the variant; Middle: number of hosts susceptible to the wild-type and the variant, with the equilibrium value $\delta/\alpha$ as a gray line; Right: fraction of the infections due to the variant. The thick gray line shows the expected equilibrium frequency $\beta$ in the case with one immune group, given in Eq. 6. The dashed line shows the trajectory of a constant fitness logistic growth with the same initial growth rate.
@@
where $\varepsilon\ll 1$. Our reasoning is that we expect an adaptive variant to escape the existing immunity for part of the host population, here immune group $1$, while having little effect on the rest of the hosts. As a result, a larger fraction of the hosts will initially be immune to the variant. If mixing is rapid, the initial growth rate of the variant is smaller by a factor $M$ compared to the one-group case. For $\varepsilon=0$, the equilibrium frequency does not depend on $M$ and Equation 6 remains valid, while for small $\varepsilon$ the equilibrium shifts slightly. If mixing is slow, the initial growth of the variant is rapid as in the one-group case, but then spreads only slowly across groups.
The central and bottom rows of Figure 1 show the result of the invasion for $M$ groups respectively for the rapid and slow mixing cases. In both scenarios, the initial number of hosts susceptible to the variant is now closer to $\delta/\alpha$, and accordingly the initial growth rate of $I^{v}$ is smaller. When mixing is fast, the frequency of the variant initially resembles a standard selective sweep (dashed line in Figure 1) before saturating, while dynamics are more complicated for the slow mixing case. Either way, the main effect of the immune groups is that the overshoot past the equilibrium is much smaller and dampening of the oscillations stronger. As a result, the frequency of the variant approaches its equilibrium value without effectively sweeping to fixation before. While this simple two-strain model predicts that the two strains come to an equilibrium at frequency $\beta$, their frequency will of course continue to change due to emergence of additional strains, which we will discuss below.
@@
If the initial growth is due to higher susceptibility, it is misleading to think of it as an intrinsic fitness advantage of the variant. What allows the initial growth is better described as an imbalance in the immune state of the host population, which gradually disappears as the variant becomes more frequent, as shown in the central panels of Figure 1. In this sense, our model is comparable to ecological systems where interaction between organisms cannot be fully explained using a scalar fitness Tikhonov and Monasson (2018). An important implication is that predicting the equilibrium frequency reached by the variant and its ultimate fate is hard from the observation of a partial frequency trajectory.
@@
In the invasion scenario discussed above, dynamics stop after the initial variant has reached an equilibrium frequency. However, as the viral population evolves, new adaptive mutants can appear. In the framework of the SIR model, a new strain translates into extending the cross-immunity matrix by one row and one column. Each new variant will perform its own partial sweep, and saturate at frequencies $\beta_{2},\beta_{3},\ldots$ This process is shown in panel A of Figure 2, using the SIR model to simulate up to $N=7$ variants. For the sake of illustration, it shows a simple scenario where the initial variant appears at time 00 in a homogeneous wild-type population, and subsequent mutants appear at regular time intervals. Simulations are performed using $M=10$ immune groups, resulting in a slight overshoot of the equilibrium frequency for each trajectory.
Figure 2: A: Simulation of SIR Equations 1 & 2 with additional strains appearing at regular time intervals. The fraction of infections (frequency) caused by each strain is shown as a function of time. The first strain to appear at $t=0$ is the variant of interest, and curves are shown in shades of red if they appear on the background of this variant, and of blue if they appear on the background of the wild-type. B: Same as A but with frequencies stacked vertically. The black line delimiting the red and blue areas represents the frequency at which the mutations defining the original variant are found. C: Three realizations of the random walk of Equation 8, all starting at $x\simeq 0.5$. Two instances converge rapidly to frequency 00 and $1$, corresponding to apparent selective sweeps, while the remaining one oscillates for a longer time. D: Representation of a partial sweep using the scalar fitness parametrization of Equation 10. The frequency $x$ of the variant is shown as a blue line saturating at value $\beta$ (gray line). The thin dashed line shows a selective sweep with constant fitness advantage $s_{0}$. The fitness $s$ is a red dashed line, using the right-axis.
@@
Examples of trajectories from the random walk are shown on panel C of Figure 2, all initially starting around $x_{0}\simeq 1/2$. Two trajectories are seen converging monotonously to 00 and $1$. This is a consequence of one interesting property of Equation 8: the probability for $\Delta x$ to be positive increases with $x$ but the magnitude of the upwards steps decreases as $1-x$, and symmetrically with downward steps. This leads to trajectories leading almost exponentially to 00 and $1$: it can in fact be shown that trajectories that always go downwards or upwards represent a finite and relatively large fraction of all possible trajectories (see SI). On the other hand, steps away from the closest boundary are unlikely but much larger, resulting in ‘jack-pot’ events (Hallatschek, 2018). This can be seen for the blue trajectory in Figure 2, which oscillates for a longer time.
@@
On the other hand, the second moment closely reminds of the theory of neutral drift Kimura (1964): in neutral evolution, allele frequency also undergoes a zero-average random walk with the second moment having the form $x(1-x)/N$ with $N$ being the population size. Therefore, this model would predict an “effective population size” as $N_{e}^{-1}=\rho\langle\beta^{2}\rangle_{P_{\beta}}$ completely independent of the size of the viral population. However, there are important difference to neutral drift: in neutral evolution, higher moments of order $k>2$ decay as $N^{1-k}$ and are thus negligible in large populations, whereas here they are independent of $N$ and scale as $\langle\beta^{k}\rangle_{P_{\beta}}$. Depending on higher moments of $P_{\beta}$, allele dynamics will deviate qualitatively from neutral behavior.
@@
In general, the dynamics of the SIR model proposed in Equations 1&2 depend on the interactions between $N$ strains through an $N\times N$ cross-immunity matrix. While this model is useful to give a mechanistic explanation of partial sweeps, it is in general impractical to analyze and numerically simulate for a large number of variants. For this reason, we propose a simpler empirical model for partial sweeps where the growth rate $s$ of the variant is not explicitly set by the susceptibility dynamics in the host population, but instead decays at a rate proportional to the frequency $x$ of the variant:.
@@
The dynamic of $x$ in the first equation is simply given by the usual logistic growth with fitness $s$. To mimic increasing immunity against the invading variant, the growth advantage $s$ decreases proportionally to the abundance of the variant (second part of Eq. 10).
@@
This final value $\beta$ depends only on the ratio between the initial fitness advantage $s_{0}$ and the rate of fitness decay $\nu$. For a large enough $s_{0}$, $\beta$ can be arbitrarily close to $1$, meaning that this model still accommodates for full selective sweeps as a special case. In the general case, $x$ reaches its final value $\beta<1$ and remains there forever unless other variants appear. Due to its simplicity, this model does not show the oscillatory behavior of the SIR model, but otherwise recapitulates the salient feature of invading immune evasive variants: (i) initial exponential growth, and (ii) eventual saturation at an intermediate frequency. This simplified model of partial sweeps leads to the same random walk dynamics as the SIR model above and allows us to study the implications of this dynamics systematically.
@@
In Figure 3, we reproduce some of our results of Barrat-Charlaix et al. (2021) and extend them to SARS-CoV-2. To quantify predictability, we ask the following question: given the state of a viral population at times $0,1,\ldots,t$, what can we say about variant frequencies at times $t+1,\ldots$? We performed a retrospective analysis of viral evolution and identified all amino-acid mutations that were observed to grow from frequency 00 to an arbitrary threshold $x^{\star}$. Adaptive beneficial mutations should in principle be overrepresented in this group. Looking at the corresponding frequency trajectories, we can see on average whether they keep on growing beyond $x^{\star}$. Figure 3 shows these trajectories for the amino acid substitutions in the HA protein of A/H3N2 influenza, using data from 2000 to 2023, and the SARS-CoV-2 genome using data from 2020 to 2023. Panels on the left show all trajectories that reached $x^{\star}=0.4$, with their average displayed in black. The panels on the right show the average trajectory for different threshold values $x^{\star}$ between $0.1$ and $0.8$.

### v3

@@ Abstract
The traveling wave framework has been extensively used in this context as it allows for a straightforward way to approach the prediction problem: each variant is assumed to have a fixed fitness relative to other variants, and inferring the fitness of all competing variants should allow prediction of the future composition of the population. Indeed, current methods typically infer the instantaneous growth advantage of a strain based on past and present circulation and then project this growth advantage forward in time Luksza and Lässig (2014); Neher et al. (2014); Huddleston et al. (2020). While future mutations can reshuffle the relative fitness of lineages and thereby limit predictability, in these models a lineage that is most fit at any given time is most likely to dominate in the long run.
@@
One short-coming of the traveling wave approach is the lack of explicit representation of the epidemiological dynamics and of the host’s immunity. Indeed, fitness is only an effective parameter that summarizes the complex interplay between viral antigenic properties and the hosts’ immune systems. As such, it cannot explicitly describe important phenomena such as the build-up of host immunity to new variants, variant specific immunity, or the interaction between strains through antigenic cross-reactivity. Taking the hosts’ immunity and viral cross-immunity into account has the potential to strongly improve predictions Meijers et al. (2023) or explain why prediction is difficult (Barrat-Charlaix et al., 2021).
@@
The interaction between epidemiological dynamics and hosts’ immunity are often modeled using generalizations of the Susceptible-Infected-Recovered model (SIR) to include multiple viral strains Gupta et al. (1998); Gog and Grenfell (2002). In this setting, the natural question is that of the ultimate fate of the pathogen: will it go extinct, diversify to the point of speciation, or reach the so-called Red Queen State where new strains continuously replace old ones Yan et al. (2019); Marchi et al. (2021); Chardès et al. (2023); Rouzine and Rozhnova (2018). To remain tractable, these studies typically approximate population immunity as a low-dimensional landscape in which the viral population evolves and ignore the complex heterogeneity in the immunity of different individuals. Furthermore, immunity is often assumed to be long-lived and evolution of the pathogen in a stable low dimensional landscape gives rise to traveling waves.
@@
Here, we study how novel variants of a virus shape the host population’s immunity, which in turn changes their own growth dynamics. To do so, we use a multi-strain SIR model combining immune waning and heterogeneous immunity of the hosts. Such heterogeneity has been demonstrated for influenza virus in individuals of different ages Lee et al. (2019); Welsh et al. (2023). We show that this model generically leads to a situation where novel immune evasive variants emerge. In a homogeneous population of hosts, this leads to a succession of selective sweeps where novel variants compete against each other and replace previously circulating variants. However, in a heterogeneous population with more rapid waning of immunity, initially growing variants lose their selective advantage before reaching fixation due to immunological adjustment of the host population. The phenomenology of our epidemiological model is reminiscent of ecological systems such as consumer resource models, where adaptation by one species shifts the global equilibrium and distribution of other species but does not necessarily result in a selective sweep Good et al. (2018). In these systems, adaptation can usually not be modeled by a fixed fitness parameter for each strain but rather depends on the composition of the population Tikhonov and Monasson (2018).
@@
Strain dynamics in our model differ qualitatively from what is expected in the traveling wave scenario. While adaptive mutations are highly overrepresented in genetic diversity, they cease having a growth advantage when reaching intermediate frequencies, a process we call “expiring fitness”. Once the fitness effect of a mutation has expired, its frequency randomly changes up or down as subsequent adaptive mutations occur on the same or on different genomic backgrounds.
@@
We describe the interaction of several viral strains and host immunity using a Susceptible/Infected compartmental model, similar to those used in Gog and Grenfell (2002); Yan et al. (2019). In the most general form, the model describes $N$ variants of the virus labeled $a\in\{1\ldots N\}$ circulating among $M$ groups of hosts with distinct exposure histories labeled $i\in\{1\ldots M\}$ (immune groups). These groups could be different age cohorts or could be geographically separated. For each group $i$, we define compartments $I_{i}^{a}$ and $S_{i}^{a}$ as respectively the number of individuals of this group infected or susceptible to strain $a$. We assume that the total population of hosts is $1$ so that we always have $0\leq I^{a}_{i},S^{a}_{i}\leq 1$, and values of $I^{a}_{i}$ and $S^{a}_{i}$ can be interpreted as fractions of the host population.
@@
When a host of group $i$ is infected by strain $b$, it gains immunity against the infecting strain $b$, but also to other strains $a$ with probability $0\leq K_{i}^{ab}\leq 1$. Thus, $S_{i}^{a}$ decreases at a rate proportional to $K_{i}^{ab}$ and to the number of hosts infected by $b$ for every strain $b$. Since susceptibles to $a$ are depleted by infections from other strains, the dynamics of all strains are coupled. This coupling is determined by the matrices $\mathbf{K}_{i}$ of dimension $N\times N$, which in general differ between groups $i$ with different prior exposure history. Additionally, waning of immunity at a rate $\gamma$ causes immune hosts to re-enter the susceptible compartment. We can now write the dynamics of $S_{i}^{a}$ as
@@
An important property of our model is that the probability to generate cross-immunity can differ between groups. The motivation is that strains $a$ and $b$ may be perceived as antigenically different by some immune systems, leading to a low $K_{i}^{ab}$, but as highly similar by others, leading to $K_{i}^{ab}\simeq 1$. Such a heterogeneous response by different immune systems has been observed experimentally in the case of influenza: in Lee et al. (2019); Welsh et al. (2023) for instance, it was found that a given mutation in an influenza strain may allow escape from the antibodies of some individuals, i.e. low $K_{i}^{ab}$, while it had little effect for the serum of other individuals, i.e. high $K_{i}^{ab}$. Heterogeneous immune response could be caused by varying histories of strain exposure for different individuals, for instance due to difference in age or geographical region. If immune groups correspond to age cohorts, mixing between groups is rapid, and we can simplify the connectivity between groups to $C_{ij}=1/M$. If immune groups are shaped by geographic differences in exposure, the connectivity would be close to 1 on the diagonal ($1-C_{ii}\ll 1$) while off-diagonal terms would be small ($C_{ij}\ll 1$ for $i\neq j$).
@@
In the first version of this scenario, we will only consider one immune group, that is $M=1$. We can thus skip the indices $i,j\in\{1\ldots M\}$, and we only have one cross-immunity matrix $\mathbf{K}$ that we parametrize as
@@
$\begin{split}\dot{S}^{a}&=-\alpha S^{a}\sum_{b\in\{wt,v\}}K^{ab}I^{b}+\gamma(1-S^{a}),\\ \dot{I}^{a}&=(\alpha S^{a}-\delta)I^{a}.\\ \end{split}$ (4)
We can immediately derive the equilibrium state for this simplified case. We first define the reproductive number of strain $a$ as $R^{a}=\alpha S^{a}/\delta$: $I^{a}$ grows when $R^{a}>1$ and declines when $R^{a}<1$. The equilibrium susceptibility is therefore $S^{a}=\delta/\alpha$, such that $R^{a}=1$. On the other hand, the equilibrium prevalence is determined by the inverse of the cross-immunity matrix $\mathbf{K}$:
@@
We are primarily interested in an “invasion” scenario where only the wild-type is initially present in the population, that is $I^{v}=0$ at $t<0$. Cross-immunity with the resident strain reduces the fraction of hosts susceptible to the variant below one even though it has not circulated yet. But the number of susceptible hosts is always larger than the equilibrium value $\delta/\alpha$ unless $f=1$, As a result, the growth rate of the variant is initially positive and given by
$s(t=0)=\frac{(1-f)(\alpha-\delta)}{\delta+f(\alpha-\delta)}$ (7)
The variant thus increases initially exponentially until it has become sufficiently frequent that it starts having a substantial effect on the immunity landscape, before eventually settling into an equilibrium with the wild-type. The details of the equilibrium reached by the system in the absence of additional mutant variants is given in the SI. Figure 1 explores different scenarios numerically.
Figure 1: Invasion scenario for Top row: one immune group, Middle row: $M=10$ immune groups and fast mixing $C_{ij}=1/M$ and Bottom row $M=10$ immune groups and slow mixing $C_{ij}=1/10M$. Other parameters are the same for all rows: in units of $\delta$, we set $\alpha=3$ and $\gamma=5\cdot 10^{-3}$, and $f=0.65$, $b=0.8$, $\varepsilon=0.01$. For both rows, graphs represent: Left: number of hosts infectious with the wild-type and the variant; Middle: number of hosts susceptible to the wild-type and the variant, with the equilibrium value $\delta/\alpha$ as a gray line; Right: fraction of the infections due to the variant. The thick gray line shows the expected equilibrium frequency $\beta$ in the case with one immune group, given in Eq. 6. The dashed line shows the trajectory of a constant fitness logistic growth with the same initial growth rate.
@@
where $\varepsilon\ll 1$. Our reasoning is that we expect an adaptive variant to escape the existing immunity for part of the host population, here immune group $1$, while having little effect on the rest of the hosts.
One consequence of many groups that are indifferent to the variant is that globally the excess susceptibility to the variant is lower. If mixing is rapid, the initial growth rate of the variant is smaller by a factor $M$ compared to the one-group case. If mixing is slow, the initial growth of the variant is as rapid as in the one-group case, but then spreads only slowly across groups. Globally, the frequency of the variant thus never reaches values close to one and population wide oscillations are reduced.
The central and bottom rows of Figure 1 show the result of the invasion for $M$ groups respectively for the rapid and slow mixing cases. In both scenarios, the initial number of hosts susceptible to the variant is now closer to $\delta/\alpha$. When mixing is fast, the frequency of the variant initially resembles a standard selective sweep (dashed line in Figure 1) before saturating, while dynamics are more complicated for the slow mixing case. Either way, the main effect of the immune groups is that the overshoot past the equilibrium is much smaller and dampening of the oscillations stronger. As a result, the frequency of the variant approaches its equilibrium value without effectively sweeping to fixation before.
Notably, the equilibrium frequency in the above examples does not depend on $M$ and Equation 6 remains valid for $\varepsilon=0$. This invariance is a consequence of the fact that for $\varepsilon=0$, the variant and wild-type strains are completely equivalent in immune groups $i>1$ and equilibrium is only determined by cross-immunity in group $i=1$ (see SI). For small $\varepsilon$ the equilibrium shifts slightly, but Equation 6 remains a good approximation.
While this simple two-strain model predicts that the two strains come to an equilibrium at frequency $\beta$, their frequency will of course continue to change due to emergence of additional strains, which we will discuss below.
@@
If the initial growth is due to higher susceptibility, it is misleading to think of it as an intrinsic fitness advantage of the variant. Instead, the initial growth is the result of an imbalance in the immune state of the host population, which gradually disappears as the variant becomes more frequent, as shown in the central panels of Figure 1. In this sense, our model is comparable to ecological systems where interaction between organisms cannot be fully explained using a fixed scalar fitness for each strain but rather depends on the composition of the viral and host population. In particular, the stalling of frequency increase giving rise to the partial sweep is reminiscent of consumer resource models Tikhonov and Monasson (2018); Good et al. (2018), highlighting the link between ecological and epidemiological models. An important consequence of these dynamics is that predicting the equilibrium frequency reached by the variant and its ultimate fate is hard from the observation of a partial frequency trajectory.
@@
In the invasion scenario discussed above, dynamics stop after the initial variant has reached an equilibrium frequency. However, as the viral population evolves, new adaptive mutants can appear. In the framework of the SIR model, a new strain translates into extending the cross-immunity matrix by one row and one column. Each new variant will perform its own partial sweep, and saturate at frequencies $\beta_{2},\beta_{3},\ldots$ sampled from some distribution $P_{\beta}$. This process is shown in panel A of Figure 2, using the SIR model to simulate up to $N=7$ variants. For the sake of illustration, it shows a simple scenario where the initial variant appears at time 00 in a homogeneous wild-type population, and subsequent mutants appear at regular time intervals. Simulations are performed using $M=10$ immune groups, resulting in a slight overshoot of the equilibrium frequency for each trajectory.
Figure 2: A: Simulation of SIR Equations 1 & 2 with additional strains appearing at regular time intervals. The fraction of infections (frequency) caused by each strain is shown as a function of time. The first strain to appear at $t=0$ is the variant of interest, and curves are shown in shades of red if they appear on the background of this variant, and of blue if they appear on the background of the wild-type. B: Same as A but with frequencies stacked vertically. The black line delimiting the red and blue areas represents the frequency at which the mutations defining the original variant are found. C: Three realizations of the random walk of Equation 9, all starting at $x\simeq 0.5$. Two instances converge rapidly to frequency 00 and $1$, corresponding to apparent selective sweeps, while the remaining one oscillates for a longer time. D: Representation of a partial sweep using the expiring fitness parametrization of Equation 11. The frequency $x$ of the variant is shown as a blue line saturating at value $\beta$ (gray line). The thin dashed line shows a selective sweep with constant fitness advantage $s_{0}$. The fitness $s$ is a red dashed line, using the right-axis.
@@
Examples of trajectories from the random walk are shown on panel C of Figure 2, all initially starting around $x_{0}\simeq 1/2$. Two trajectories converge monotonically to 00 and $1$. This is a consequence of one interesting property of Equation 9: the probability for $\Delta x$ to be positive increases with $x$, but the magnitude of the upwards steps decreases as $1-x$, and symmetrically with downward steps. This leads to trajectories leading almost exponentially to 00 and $1$: it can in fact be shown that trajectories that always go downwards or upwards represent a finite and relatively large fraction of all possible trajectories (see SI). On the other hand, steps away from the closest boundary are unlikely but much larger, resulting in ‘jack-pot’ events (Hallatschek, 2018). This can be seen for the blue trajectory in Figure 2, which oscillates for a longer time.
@@
On the other hand, the second moment resembles neutral drift Kimura (1964): in neutral evolution, allele frequency also undergoes a zero-average random walk with the second moment having the form $x(1-x)/N$ with $N$ being the population size. Therefore, this model would predict an “effective population size” as $N_{e}^{-1}=\rho\langle\beta^{2}\rangle_{P_{\beta}}$ completely independent of the size of the viral population. However, there are important differences to neutral drift: in neutral evolution, higher moments of order $k>2$ decay as $N^{1-k}$ and are thus negligible in large populations, whereas here they are independent of $N$ and scale as $\langle\beta^{k}\rangle_{P_{\beta}}$. Depending on higher moments of $P_{\beta}$, allele dynamics will deviate qualitatively from neutral behavior.
@@
In general, the dynamics of the SIR model proposed in Equations 1&2 depend on the interactions between $N$ strains through an $N\times N$ cross-immunity matrix. While this model is useful to give a mechanistic explanation of partial sweeps, it is in general impractical to analyze and numerically simulate for a large number of variants. The random walk model introduced above is simple to analyze and simulate, but assumes that variants rise to their equilibrium frequency instantaneously.
To explore the consequences of partial sweeps over broader parameter ranges, we propose an empirical model that has the same qualitative properties as the overdamped SIR, namely a growth rate that decreases as a strain becomes more frequent and partial sweep trajectories, but is simpler to analyze and simulate on a large scale. In this effective model, the growth rate $s$ of the variant is not explicitly set by the susceptibility dynamics in the host population, but instead decays at a rate proportional to the frequency $x$ of the variant:
@@
The dynamic of $x$ in the first equation is simply given by the usual logistic growth with fitness $s$. To mimic increasing immunity against the invading variant, the growth advantage $s$ decreases proportionally to the abundance of the variant (second part of Eq. 11). The initial value of $s_{0}$ is connected to the invasion rate of the SIR models given in Eq. 38 of the SI.
@@
This final value $\beta$ depends only on the ratio between the initial fitness advantage $s_{0}$ and the rate of fitness decay $\nu$. For a large enough $s_{0}$, $\beta$ can be arbitrarily close to $1$, meaning that this model still accommodates for full selective sweeps as a special case. In the general case, $x$ reaches its final value $\beta<1$ and remains there forever unless other variants appear.
It is important to state that the main aim of this effective model is to qualitatively reproduce the phenomenology of the SIR, and in particular the partial sweeps, while being easier to simulate. It recapitulates the salient feature of invading immune evasive variants: (i) initial exponential growth, and (ii) eventual saturation at an intermediate frequency. We can thus use it to analyze the long term consequences of the random walk dynamics of Figure 2. However, we do not expect the frequency of the variant $x$ to have quantitatively equivalent dynamics in the two models. In particular, due to its simplicity, this model does not show the complex oscillatory behavior of the SIR model. Section (A.9) of the SI discusses in more detail the links between the parameters of the two models and the fundamental differences. While we can express the rate $\nu$ at which the growth rate declines in terms of the parameters of the simplest SIR models, for models with many groups or with oscillatory dynamics, the decay rate of the growth advantage should be interpreted as an effective parameter that captures a generic effect of reduced growth with increasing circulation.
@@
In Figure 3, we reproduce some of our results of Barrat-Charlaix et al. (2021) and extend them to SARS-CoV-2. To quantify predictability, we ask the following question: given the state of a viral population at times $0,1,\ldots,t$, what can we say about variant frequencies at times $t+1,\ldots$? We performed a retrospective analysis of viral evolution and identified all amino-acid mutations that were observed to grow from frequency 00 to an arbitrary threshold $x^{\star}$. Adaptive beneficial mutations should in principle be overrepresented in this group and if they provide a persistent fitness advantage, we would expect them on average to keep on growing beyond $x^{\star}$. Figure 3 shows these trajectories for the amino acid substitutions in the HA protein of A/H3N2 influenza, using data from 2000 to 2023, and the SARS-CoV-2 genome using data from 2020 to 2023. Panels on the left show all trajectories that reached $x^{\star}=0.4$, with their average displayed in black. The panels on the right show the average trajectory for different threshold values $x^{\star}$ between $0.1$ and $0.8$.